### v1

# Contrastive Disentangling: Fine-grained representation learning through multi-level contrastive learning without class priors

Abstract
Recent advancements in unsupervised representation learning often leverage class information to enhance feature extraction and clustering performance. However, this reliance on class priors limits the applicability of such methods in real-world scenarios where class information is unavailable or ambiguous. In this paper, we propose Contrastive Disentangling (CD), a simple and effective framework that learns representations without any reliance on class priors. Our framework employs a multi-level contrastive learning strategy that combines instance-level and feature-level losses with a normalized entropy loss to learn semantically rich and fine-grained representations. Specifically, (1) the instance-level contrastive loss encourages the separation of feature representations for different samples, (2) the feature-level contrastive loss promotes independence among the feature head predictions, and (3) the normalized entropy loss encourages the feature heads to capture meaningful and prevalent attributes from the data. These components work together to enable CD to significantly outperform existing methods, as demonstrated by extensive experiments on benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10, particularly in scenarios where class priors are absent. The code is available at https://github.com/Hoper-J/Contrastive-Disentangling.
keywords:
Disentangling representations , Clustering , Contrastive learning , Representation learning , Unsupervised learning
[…]
[fafu]organization=School of computer science and information engineering, addressline=Fujian Agriculture and Forestry University, city=Fuzhou, postcode=350002, country=China
1 Introduction
The growing importance of unsupervised learning in processing large-scale unlabeled datasets has increased the demand for effective feature extraction techniques that do not rely on predefined class labels. In computer vision tasks, unsupervised representation learning has become a powerful solution for deriving meaningful feature representations from raw data. A key advancement in this domain is contrastive learning, which has emerged as a prominent paradigm within unsupervised representation learning. By maximizing consistency between different augmented views of the same data instance while minimizing similarity between different instances, contrastive learning has significantly advanced feature extraction (Ye et al., 2019; Chen et al., 2020).
As research in this field progresses, numerous contrastive learning methods (Van Gansbeke et al., 2020; Li et al., 2021; Qian, 2023) have started to utilize class number information to further improve performance. These approaches use class numbers as prior knowledge during network construction, achieving state-of-the-art results in unsupervised clustering tasks. This suggests that incorporating class information during training yield more discriminative feature representations. However, in practical applications, class information is often unavailable, or the data may lack clear structure, limiting the applicability of these methods.
To address this limitation, we propose a novel framework, Contrastive Disentangling (CD), which enhances feature extraction and disentanglement in unsupervised learning tasks through a multi-level contrastive learning strategy and an innovative disentanglement loss function, all without relying on class number information. Unlike methods that rely on additional complex modules, such as self-labeling, our approach focuses on validating the effectiveness of the core framework to ensure its broad applicability.
Figure 1: Multi-level Contrastive Learning in the CD Framework. Instance-level (top) operates on image instances, while feature-level (bottom) focuses on features visualized with LIME.
As illustrated in Figure 1, the CD framework incorporates two levels of contrastive learning: instance-level and feature-level. Instance-level contrastive learning aims to distinguish different image instances within the feature space, while feature-level contrastive learning seeks to differentiate the features captured by distinct prediction heads. In the feature-level module, we employ the LIME (Local Interpretable Model-Agnostic Explanations) technique (Ribeiro et al., 2016) to interpret the outputs of specific feature prediction heads, providing an intuitive visualization of the diverse feature attributes captured by the model across the image batch. In summary, the main contributions of this paper include:
  1. 1.
    We introduce feature prediction heads, a novel addition to the feature learning process. Using normalized entropy loss, these heads effectively capture common features prevalent in the dataset, thereby enhancing the representational capacity and generalization of the features.
  2. 2.
    A feature-level contrastive learning strategy is proposed, ensuring that different feature prediction heads focus on learning distinct and complementary features, thus promoting feature disentanglement.
  3. 3.
    We propose a simple and easy-to-use framework that significantly enhances the learning of fine-grained features in unsupervised learning tasks without relying on class information. Extensive experiments across multiple benchmark datasets demonstrate the superior performance of the CD framework, particularly in metrics like NMI, where it outperforms compared methods. Through comprehensive ablation studies, we validate the effectiveness of each component, providing new insights in unsupervised representation learning.
2 Related Work
2.1 Unsupervised Representation Learning
Unsupervised representation learning seeks to extract meaningful features from unlabeled data. The field is primarily dominated by two approaches: generative models and contrastive learning. Generative models (Bengio et al., 2006; Kingma, 2013; Goodfellow et al., 2014) learn to generate samples resembling the true data distribution by modeling the underlying latent space. However, these models often struggle to learn invariant feature representations, which are insensitive to certain input transformations like rotation, translation, and scaling (Ntelemis et al., 2023).
To overcome these challenges, instance-level contrastive learning methods (He et al., 2020; Chen et al., 2020; Grill et al., 2020; Chen and He, 2021) have been proposed. These approaches significantly enhance the model’s capacity to extract invariant features by maximizing the similarity between different views of the same instance. These methods have exhibited superior performance across various downstream tasks, underscoring their efficacy in unsupervised learning scenarios.
2.2 Utilization of Class Number Information
Recent approaches have leveraged class number information as additional prior knowledge to enhance model performance. Methods (Van Gansbeke et al., 2020; Li et al., 2021; Niu et al., 2022; Qian, 2023) incorporate class number to constrain the final output during network construction. Several of these methods (Van Gansbeke et al., 2020; Niu et al., 2022; Qian, 2023) also employ pseudo-labeling techniques, further enhancing the model’s performance in unsupervised clustering tasks. These advancements indicate that class-level information can effectively assist models in distilling more semantically meaningful features.
However, this dependence on class information constrains the applicability of these methods in real-world settings where class labels or the number of classes are frequently unknown or ambiguous. This limitation emphasizes the necessity for more adaptive approaches that can perform effectively without relying on class priors.
2.3 Multi-level Contrastive Learning
The introduction of cluster-level contrastive learning by CC (Li et al., 2021) for clustering tasks has unveiled new avenues in contrastive learning research. However, partitioning based on predefined categories may not adequately capture the fine-grained features inherent in datasets, potentially constraining the model’s feature extraction capabilities. Building upon this observation, we propose a novel architecture that leverages feature space representations as inputs to feature prediction heads. By extracting common features and implementing feature-level contrastive learning, our approach achieves relative decoupling of distinct feature heads. This facilitates the capture of more discriminative and fine-grained features, thereby advancing the state-of-the-art in unsupervised representation learning.
3 Method
3.1 Proposed Contrastive Disentangling Framework
The proposed CD framework leverages multi-level contrastive learning, as illustrated in Figure 2. It consists of the following key components:
Figure 2: Architecture of the proposed Contrastive Disentangling (CD) framework. The top section illustrates the shared backbone network, instance-level projector, and feature-level predictor, while the bottom section displays the optimization objectives for instance-level and feature-level contrastive learning.
  • 1.
    Data Augmentation. Each input batch $x$ undergoes a series of data augmentation operations, such as random cropping, horizontal flipping, color distortion, and blurring. These augmentations generate two correlated views, denoted as $x^{(1)}$ and $x^{(2)}$, enabling the model to focus on semantically relevant features that remain invariant under these transformations.
  • 2.
    Feature Extraction via ResNet Backbone. A ResNet architecture (He et al., 2016) is used as the shared backbone network $f(\cdot)$, which extracts feature representations from the augmented views $x^{(1)}$ and $x^{(2)}$. This choice aligns with prevailing practices in contrastive learning (Chen et al., 2020; Li et al., 2021). We denote the extracted feature representations as $h^{(1)}=f(x^{(1)})$ and $h^{(2)}=f(x^{(2)})$.
  • 3.
    Instance Projector. The feature representations $h^{(1)}$ and $h^{(2)}$ are passed through an instance projector $g(\cdot)$, implemented as a multi-layer perceptron (MLP). This MLP consists of two linear layers, batch normalization (Ioffe, 2015), and a ReLU activation function. The output is a set of latent representations $z^{(1)}=g(h^{(1)})$ and $z^{(2)}=g(h^{(2)})$, which are crucial for the contrastive learning process (Chen et al., 2020).
  • 4.
    Feature Predictor. The latent representations $z^{(1)}$ and $z^{(2)}$ are fed into a feature predictor $h(\cdot)$, structured similarly to the instance projector but using sigmoid activation in the final layer. The feature predictor outputs predictions $y^{(1)}=h(z^{(1)})$ and $y^{(2)}=h(z^{(2)})$, organized into multiple prediction heads. Each head predicts a distinct feature attribute. We will validate the design of the feature predictor in Section 4.5.3.
  • 5.
    Loss Functions. The total loss comprises three components: the instance-level contrastive loss, the feature-level contrastive loss, and an additional normalized entropy loss. These losses work together to enhance feature diversity, promote independence among learned features, and ensure a wide range of feature predictions.
3.2 Instance-level Contrastive Learning
The objective of instance-level contrastive learning is to pull the representations of different augmented views of the same image closer in feature space while pushing apart representations of different images. Specifically, given an input batch $x$, two views $x^{(1)}$ and $x^{(2)}$ are generated via data augmentation, which are then processed by the shared backbone network and instance projector to yield latent representations $z^{(1)}=g(f(x^{(1)}))$ and $z^{(2)}=g(f(x^{(2)}))$.
These latent representations, $z^{(1)}$ and $z^{(2)}$, are concatenated into a matrix $\mathbf{z}\in\mathbb{R}^{2N\times d}$, where $N$ is the batch size and $d$ is the feature dimension:
$\mathbf{z}=\begin{bmatrix}z^{(1)}\\ z^{(2)}\end{bmatrix}=\begin{bmatrix}z_{1}\\ z_{2}\\ \vdots\\ z_{2N}\end{bmatrix}$ (1)
To quantify the similarity between any two latent representations, we introduce the cosine similarity measure $\text{sim}_{z}(i,j)$:
$\text{sim}_{z}(i,j)=\frac{z_{i}\cdot z_{j}}{\|z_{i}\|\|z_{j}\|}$ (2)
where $i,j\in\{1,2,\dots,2N\}$ and $N$ is the batch size. Based on this similarity measure, we optimize the model by maximizing the similarity between different views of the same instance while minimizing the similarity between different instances. The loss function for instance $i$ is defined as:
$\ell_{i}^{\text{inst}}=-\log\frac{\exp(\text{sim}_{z}(i,\text{pos}(i))/\tau_{\text{inst}})}{\sum_{k=1}^{2N}\mathbbm{1}_{[k\neq i]}\exp(\text{sim}_{z}(i,k)/\tau_{\text{inst}})}$ (3)
Here, $\text{pos}(i)$ represents the index of the positive sample corresponding to $i$. If $i\leq N$, then $\text{pos}(i)=i+N$; otherwise, $\text{pos}(i)=i-N$. The temperature parameter $\tau_{\text{inst}}$ within the instance-level contrastive loss governs the smoothness of the similarity distribution. Additionally, the indicator function $\mathbbm{1}_{[k\neq i]}$ assigns a value of 1 when $k\neq i$, thereby excluding the self-similarity of $z_{i}$. The instance-level contrastive loss is then calculated across all instance views:
$\mathcal{L}_{\text{inst}}=\frac{1}{2N}\sum_{i=1}^{2N}\ell_{i}^{\text{inst}}$ (4)
3.3 Disentanglement Loss
To enhance the independence and diversity of the learned features, we introduce two complementary loss functions: feature-level contrastive loss and normalized entropy loss. These losses work together to promote the learning of disentangled and fine-grained feature attributes.
Feature-level Contrastive Loss. The feature predictor $h(\cdot)$ is designed with multiple independent prediction heads, each focused on predicting a fine-grained feature attribute. After processing the input views $x^{(1)}$ and $x^{(2)}$ through the backbone network and instance projector to generate latent representations $z^{(1)}$ and $z^{(2)}$, these representations are fed into the feature predictor $h(\cdot)$, yielding the corresponding feature predictions $y^{(1)}$ and $y^{(2)}$.
To align with the instance-level contrastive loss, these predictions $y$ are transposed and organized into a matrix $\mathbf{y}\in\mathbb{R}^{2K\times N}$, where $K$ denotes the number of features and $N$ represents the batch size:
$\mathbf{y}=\begin{bmatrix}y^{(1)\top}\\ y^{(2)\top}\end{bmatrix}=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{2K}\end{bmatrix}$ (5)
In this matrix, $y_{i}$ and $y_{\text{pos}(i)}$ correspond to the prediction results of feature head $i$ on the same batch of samples under two different views. We then define the similarity function $\text{sim}_{y}(i,j)$ to measure the cosine similarity between any two feature prediction vectors:
$\text{sim}_{y}(i,j)=\frac{y_{i}\cdot y_{j}}{\|y_{i}\|\|y_{j}\|}$ (6)
To ensure that feature head $i$ learns features different from other heads, the feature-level contrastive loss is defined as follows:
$\ell_{i}^{\text{feat}}=-\log\frac{\exp(\text{sim}_{y}(i,\text{pos}(i))/\tau_{\text{feat}})}{\sum_{k=1}^{2K}\mathbbm{1}_{[k\neq i]}\exp(\text{sim}_{y}(i,k)/\tau_{\text{feat}})}$ (7)
Here, $\tau_{\text{feat}}$ is the temperature parameter for the feature-level contrastive loss, which controls the degree of similarity enforced between different features. The primary objective is to encourage each prediction head to learn disentangled feature attributes. The final feature-level contrastive loss is calculated as follows:
$\mathcal{L}_{\text{feat}}=\frac{1}{2K}\sum_{i=1}^{2K}\ell_{i}^{\text{feat}}$ (8)
Normalized Entropy Loss. To further promote the learning of common features, we introduce a normalized entropy loss $\mathcal{L}_{\text{entropy}}$. This loss encourages the model to produce balanced and varied predictions by maximizing the entropy of the predicted outputs. By doing so, it prevents the model from overfitting to specific features and promotes the capture of more generalizable attributes. The normalized binary entropy $H(y_{i})$ for each feature prediction vector $y_{i}$ is defined as:
$H(y_{i})=-\frac{1}{N\log(2)}\sum_{j=1}^{N}\left(y_{ij}\log(y_{ij})+(1-y_{ij})\log(1-y_{ij})\right)$ (9)
Subsequently, the normalized entropy loss $\mathcal{L}_{\text{entropy}}$ is computed as the mean entropy across all feature predictions:
$\mathcal{L}_{\text{entropy}}=\frac{1}{2K}\sum_{i=1}^{2K}H(y_{i})$ (10)
Finally, the disentanglement loss $\mathcal{L}_{\text{disentanglement}}$ combines the feature-level contrastive loss and the normalized entropy loss:
$\mathcal{L}_{\text{disentanglement}}=\mathcal{L}_{\text{feat}}-\alpha\cdot\mathcal{L}_{\text{entropy}}$ (11)
where $\alpha$ is a hyperparameter that controls the weight of the normalized entropy loss. It balances the learning of common features within the dataset and the disentanglement of independent feature representations. Minimizing this combined loss enables our CD framework to effectively facilitate the unsupervised extraction and disentanglement of meaningful features from the input data. The efficacy of $\mathcal{L}_{\text{disentanglement}}$ is further validated in the experimental section 4.5.3.
Input: Dataset $X$, Batch size $N$, Feature number $K$, Temperature constants $\tau_{\text{inst}}$ and $\tau_{\text{feat}}$
Output: Total loss $\mathcal{L}_{\text{total}}$
for each minibatch $\mathbf{x}$ of size $N$ in $X$ do
      // Apply data augmentation
       $\mathbf{x}\rightarrow\{\mathbf{x}^{(1)},\mathbf{x}^{(2)}\}$;
      for each view $\mathbf{x}^{(v)},v\in\{1,2\}$ do
             // Extract features
             $\mathbf{h}^{(v)}=f(\mathbf{x}^{(v)})$;
            // Project to latent space
             $\mathbf{z}^{(v)}=g(\mathbf{h}^{(v)})$;
            // Generate feature head predictions
             $\mathbf{y}^{(v)}=h(\mathbf{z}^{(v)})$;
       end for
      Combine $z^{(1)}$, $z^{(2)}$ by Eq.(1) and $y^{(1)}$, $y^{(2)}$ by Eq.(5);
      Compute cosine similarities $\text{sim}_{z}(i,j)$, $\text{sim}_{y}(i,j)$ by Eq. (2), (6);
      Calculate instance-level contrastive loss $\mathcal{L}_{\text{inst}}$ by Eq. (3), (4);
      Calculate feature-level contrastive loss $\mathcal{L}_{\text{feat}}$ by Eq. (7), (8);
      Calculate normalized entropy loss $\mathcal{L}_{\text{entropy}}$ by Eq. (9), (10);
      Combine all losses $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{inst}}+\mathcal{L}_{\text{feat}}-\alpha\cdot\mathcal{L}_{\text{entropy}}$;
      Update parameters of $f$, $g$, and $h$ to minimize $\mathcal{L}_{\text{total}}$;
end for
return $\mathcal{L}_{\text{total}}$
Algorithm 1 Contrastive Disentangling Framework
4 Experiment
4.1 Datasets
We conducted our experiments on four widely-used computer vision datasets: CIFAR-10, CIFAR-100 (Krizhevsky and Hinton, 2009), STL-10 (Coates et al., 2011), and ImageNet-10 (Deng et al., 2009). These datasets serve as benchmarks to rigorously evaluate the effectiveness of our proposed Contrastive Disentangling (CD) framework. Table 1 provides an overview of the dataset splits, sample sizes, and the number of classes used.
Table 1: Overview of datasets used for training and evaluation
Dataset Split Samples Classes
CIFAR-10 train+test 60,000 10
CIFAR-100-20 train+test 60,000 20
STL-10 train+test 13,000 10
ImageNet-10 train 13,000 10
For the CIFAR-100 dataset, we used 20 super-classes as ground-truth labels instead of the 100 fine-grained classes, aligning with benchmark settings in previous studies. Furthermore, to ensure the rigor of our architecture evaluation, we did not use any unseen data during training, such as the unlabeled data in STL-10.
4.2 Experimental Settings
Data Augmentation. To ensure fair comparison with existing methods (Li et al., 2021), we adopted the same data augmentation strategies and used ResNet-34 as the backbone network for feature extraction. Specifically, all datasets underwent random cropping to 224x224, random horizontal flipping, color jittering, and grayscale transformation. Notably, Gaussian blur was applied only to the ImageNet-10 dataset, as upscaling images to 224x224 for smaller datasets (e.g., CIFAR-10, CIFAR-100, and STL-10) inherently introduces a similar blurring effect.
Training Details. In our experiments, the hidden dimension and the number of features were both set equal to the batch size. We employed a cosine decay schedule without restarts (Chen et al., 2020; Loshchilov and Hutter, 2016) and gradient clipping to enhance the stability of model training. The effectiveness of these optimizations is demonstrated in the ablation study presented in Table 2. Additionally, no other techniques (e.g., self-labeling) were used to boost model performance. The hyperparameter $\alpha$, which balances the diversity of feature representations, was simply set to 1. The remaining training settings were consistent with prior contrastive learning work (Li et al., 2021): a random seed of 42 was used, and training was conducted for 1000 epochs. The Adam optimizer (Kingma, 2014) was employed, with a learning rate of $3\times 10^{-4}$ and no weight decay. The temperature parameters $\tau_{\text{inst}}$ and $\tau_{\text{feat}}$ were set to 0.5 and 1, respectively.
Environment. The experiments were conducted on an NVIDIA RTX 3090 GPU (24GB) and an 18 vCPU AMD EPYC 9754 128-Core Processor. We experimented with batch sizes of 128 and 256. The training time for the CIFAR-10 and CIFAR-100 datasets was approximately 48 GPU hours, while for the STL-10 and ImageNet-10 datasets, the training time was around 11 GPU hours. To ensure reproducibility and transparency, detailed logs of all experimental procedures are available at here.
Evaluation Metrics. We employed three clustering metrics — Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), and Accuracy (ACC) — to evaluate the feature extraction and disentanglement capabilities of the model.
4.3 Compared Methods
To rigorously assess the performance of our proposed method in various unsupervised learning scenarios, we compared it against a diverse set of representative methods, chosen for their relevance to the challenges addressed by our approach. Alongside the traditional K-means (MacQueen et al., 1967) clustering algorithm, we evaluated several fully unsupervised models—AE (Bengio et al., 2006), DeCNN (Zeiler et al., 2010), DAE (Vincent et al., 2010), VAE (Kingma, 2013), and DCGAN (Radford, 2015)—that span a spectrum of approaches from classical techniques to modern deep generative models. These methods, like our proposed Contrastive Disentangling (CD) model, do not rely on any prior class information during training, including the number of classes.
We also compared our method with unsupervised models that incorporate class number information—such as JULE (Yang et al., 2016), DEC (Xie et al., 2016), DAC (Chang et al., 2017), DCCM (Wu et al., 2019), PICA (Huang et al., 2020), and CC (Li et al., 2021)—to illustrate how our approach performs relative to methods that leverage this additional knowledge. To ensure a fair and consistent comparison, all experiments were conducted using the full set of samples, adhering to identical training and evaluation protocols.
4.4 Experimental Results
We evaluated our model at two key stages: feature extraction and feature disentanglement. Specifically, we applied K-means clustering to both the backbone network output and the final model output to assess the model’s capabilities in feature extraction and disentanglement, respectively.
Feature Extraction. Table 2 presents a comparative analysis of various models at the feature extraction stage, utilizing the NMI, ARI, and ACC metrics. Across four datasets—CIFAR-10, CIFAR-100-20, STL-10, and ImageNet-10—our Contrastive Disentanglement (CD) method, implemented with batch sizes of 128 (CD-128) and 256 (CD-256), demonstrated superior performance compared to other fully unsupervised learning models, including AE, DeCNN, DAE, VAE, and DCGAN. In comparison to the contrastive learning method CC, which utilizes prior information about the number of classes, CD consistently produced the highest NMI scores across all datasets, with notable improvements in STL-10 and ImageNet-10, while maintaining superior overall performance in most other cases.
Table 2: Feature extraction performance comparison across models. The 1st/2nd best results are indicated in bold/underlined
Dataset Metrics K-means AE DeCNN DAE VAE DCGAN CC CD-128 CD-256
[…]
Feature Disentanglement. Table 3 presents a comparison of our method with other unsupervised methods that utilize number of classes information at the final output stage. The results demonstrate that the CD method exhibited excellent performance across all metrics, particularly on the STL-10 and ImageNet-10 datasets. For CIFAR-10 and CIFAR-100-20, although the CD method’s ARI and ACC metrics were slightly lower than those of the CC method, this may be attributed to the CD method not utilizing the number of classes information to constrain the final feature prediction. Notably, on the STL-10 dataset, while the PICA method achieved a slightly higher ACC (0.713) due to the use of additional unlabeled data, the CD-128 model reached a comparable ACC of 0.702 without using extra data, and the CD-256 model further improved this to 0.734.
Table 3: Final output performance comparison across models. The 1st/2nd best results are indicated in bold/underlined, and models that utilized unlabeled datasets during training for STL-10 are marked with
Dataset Metrics JULE DEC DAC DCCM PICA CC CD-128 CD-256
[…]
These results provide strong evidence that the CD method can achieve excellent performance without requiring the number of classes information. It is particularly noteworthy that the CD method exhibited outstanding performance even with smaller batch sizes. For a detailed visualization of the CD-256 model’s feature representations using t-SNE, please see A.
4.5 Ablation Studies
To demonstrate the effectiveness of each component in our model, we conducted a series of ablation studies using the STL-10 and ImageNet-10 datasets with a batch size of 128. In sections 4.5.1 and 4.5.2, we incrementally added modules to validate their effectiveness based on the framework and objective function shown in Figure 2. In sections 4.5.3 and 4.5.4, we removed certain modules to verify their necessity.
4.5.1 Normalized Entropy Loss and Batch Normalization
The impact of normalized entropy loss (NE) and batch normalization (BN) on model performance was initially evaluated, as shown in Table 4. The experimental results demonstrate that both normalized entropy loss and batch normalization significantly enhance model performance, with their synergistic application yielding even more substantial performance improvements. Notably, NE exhibited particularly significant improvements during the feature disentanglement phase, underscoring its pivotal role in facilitating effective feature disentanglement.
Table 4: Effect of normalized entropy loss (NE) and batch normalization (BN) on performance across backbone and final output.
Dataset NE BN Backbone Feature
NMI ARI ACC NMI ARI ACC
STL-10 - - $0.519\pm 0.029$ $0.265\pm 0.037$ $0.538\pm 0.050$ $0.352\pm 0.040$ $0.065\pm 0.028$ $0.354\pm 0.041$
- $0.638\pm 0.021$ $0.482\pm 0.035$ $0.695\pm 0.031$ $0.656\pm 0.005$ $0.527\pm 0.004$ $0.671\pm 0.007$
- $0.659\pm 0.017$ $0.495\pm 0.036$ $0.653\pm 0.037$ $0.367\pm 0.052$ $0.069\pm 0.034$ $0.306\pm 0.044$
$\mathbf{0.677\pm 0.020}$ $\mathbf{0.547\pm 0.031}$ $\mathbf{0.705\pm 0.039}$ $\mathbf{0.669\pm 0.006}$ $\mathbf{0.543\pm 0.016}$ $\mathbf{0.681\pm 0.020}$
ImageNet-10 - - $0.632\pm 0.022$ $0.287\pm 0.041$ $0.620\pm 0.042$ $0.549\pm 0.038$ $0.163\pm 0.037$ $0.424\pm 0.040$
- $0.748\pm 0.020$ $0.532\pm 0.054$ $0.779\pm 0.047$ $0.887\pm 0.002$ $0.861\pm 0.002$ $0.927\pm 0.001$
- $0.837\pm 0.018$ $0.733\pm 0.050$ $0.831\pm 0.055$ $0.538\pm 0.043$ $0.152\pm 0.045$ $0.416\pm 0.055$
$\mathbf{0.857\pm 0.021}$ $\mathbf{0.772\pm 0.068}$ $\mathbf{0.858\pm 0.071}$ $\mathbf{0.890\pm 0.004}$ $\mathbf{0.864\pm 0.006}$ $\mathbf{0.929\pm 0.004}$
To further illustrate the impact of normalized entropy loss on feature disentanglement, a visual representation of the model outputs is presented in Figure 3.
(a) Default
(b) noNEloss
Figure 3: Visualization of transposed model outputs with/without normalized entropy loss (NE loss). Each row corresponds to the prediction of a specific feature attribute across the entire batch. The left visualization (a) shows the output with NE loss enabled, focusing on common features, while the right visualization (b) shows the output without NE loss, capturing more unique, image-specific features.
This visualization clearly demonstrates the effectiveness of NE loss in guiding the model to capture common features across the dataset, rather than focusing on unique features specific to individual images. The rows in Figure 3 represent the feature attributes learned by our model, as depicted in the feature-level part of Figure 1 and further visualized in B. This visual evidence corroborates the quantitative results and emphasizes the significance of NE loss in the proposed methodology.
4.5.2 Training Tricks
Following the validation of core modules, we further evaluated the contribution of cosine decay schedule and gradient clipping training techniques, as shown in Table 5.
Table 5: Effect of cosine decay schedule (Sch.) and gradient clipping (Clip.) on model performance.
Dataset Sch. & Clip. NMI ARI ACC
STL-10 - $0.669\pm 0.006$ $0.543\pm 0.016$ $0.681\pm 0.020$
$\mathbf{0.679\pm 0.007}$ $\mathbf{0.564\pm 0.023}$ $\mathbf{0.720\pm 0.026}$
ImageNet-10 - $0.890\pm 0.004$ $0.864\pm 0.006$ $0.929\pm 0.004$
$\mathbf{0.898\pm 0.003}$ $\mathbf{0.868\pm 0.006}$ $\mathbf{0.931\pm 0.003}$
The results show that both the cosine decay schedule and gradient clipping further enhance the model’s final performance and improve the stability of the results.
4.5.3 Feature Prediction Head
Next, we evaluated the effect of removing the feature prediction head, which directly reduces the model to an instance-level contrastive learning architecture, similar to SimCLR. This experiment was designed to verify the critical role of the disentanglement loss $\mathcal{L}_{\text{disentanglement}}$in the feature disentanglement process.
Table 6: Effect of feature-level head on model performance.
Dataset Feature-Level Head NMI ARI ACC
STL-10 - $0.599\pm 0.029$ $0.458\pm 0.043$ $0.649\pm 0.053$
$\mathbf{0.679\pm 0.007}$ $\mathbf{0.564\pm 0.023}$ $\mathbf{0.720\pm 0.026}$
ImageNet-10 - $0.708\pm 0.031$ $0.413\pm 0.060$ $0.608\pm 0.065$
$\mathbf{0.898\pm 0.003}$ $\mathbf{0.868\pm 0.006}$ $\mathbf{0.931\pm 0.003}$
As shown in Table 6, the results demonstrate that the feature prediction head significantly improves the model’s performance in the final layers (see Table 9 for backbone results). Removing this head leads to a noticeable degradation in performance. This confirms the critical contribution of the feature-level contrastive learning strategy, which enhances the model’s ability to disentangle features across different prediction heads.
4.5.4 Data Augmentation
This section verifies the necessity of the dual-view data augmentation strategy for feature disentanglement.
Table 7: Effect of single-view vs. dual-view data augmentation on model performance.
Dataset $x^{(1)}$ $x^{(2)}$ NMI ARI ACC
STL-10 - $0.643\pm 0.004$ $0.555\pm 0.012$ $0.717\pm 0.014$
$\mathbf{0.679\pm 0.008}$ $\mathbf{0.564\pm 0.023}$ $\mathbf{0.720\pm 0.026}$
ImageNet-10 - $0.874\pm 0.001$ $0.847\pm 0.001$ $0.918\pm 0.001$
$\mathbf{0.898\pm 0.003}$ $\mathbf{0.868\pm 0.006}$ $\mathbf{0.931\pm 0.003}$
As shown in Table 7, the results demonstrate that dual-view augmentation is crucial in the feature disentanglement stage for capturing common features in the dataset. However, in the feature extraction stage, single-view augmentation can achieve satisfactory performance (see Table 10 for details).
5 Conclusion
In this work, we proposed the Contrastive Disentangling (CD) framework, a novel approach for enhancing feature extraction and disentanglement in unsupervised learning without relying on prior knowledge of class numbers. The CD framework integrates multi-level contrastive learning with an innovative disentanglement loss function, emphasizing simplicity and core effectiveness. Experimental results demonstrate that CD consistently outperforms class-dependent methods with simple architectures (i.e., without additional structures such as self-labeling), particularly in NMI, across multiple benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10. Moreover, CD exhibits robust performance in small-batch settings, effectively capturing common features across diverse data distributions. Comprehensive ablation studies corroborate the efficacy of each framework component, validating its overall effectiveness and potential for wide-scale application in unsupervised learning scenarios.
6 Acknowledgments
This paper was supported by the Fujian Provincial Natural Science Foundation of China (No.2021J01129), the Fujian Provincial Higher Education Technology Research Association Fund Project (No.H2000134A), and the Fujian Agriculture and Forestry University Horizontal Technology Innovation Fund (No.KHF190015).
Appendix A t-SNE Visualization of Backbone and Feature Predictor Outputs
In this appendix, we present the t-SNE (t-Distributed Stochastic Neighbor Embedding) visualizations (Van der Maaten and Hinton, 2008) of the outputs both the backbone feature extraction layer and the feature predictor of our model, across all datasets: CIFAR-10, CIFAR-100, STL-10, and ImageNet-10. These visualizations are shown in Figures 4-7, where (a) represents the backbone output and (b) represents the feature predictor output for each dataset.
(a) Backbone
(b) Feature
Figure 4: t-SNE visualization for CIFAR-10
(a) Backbone
(b) Feature
Figure 5: t-SNE visualization for CIFAR-100
(a) Backbone
(b) Feature
Figure 6: t-SNE visualization for STL-10
(a) Backbone
(b) Feature
Figure 7: t-SNE visualization for ImageNet-10
Appendix B Visualizing Feature Prediction Heads with LIME
We utilize the LIME (Local Interpretable Model-Agnostic Explanations) technique (Ribeiro et al., 2016) to visualize the outputs of the feature prediction heads. LIME is configured with $num\_samples$=5000, meaning that for each image, LIME generates 5000 perturbed samples to create an interpretable explanation. We only visualize images where the feature prediction heads produce outputs greater than 0.75.
Figure 8: Visualization of the attention from five feature prediction heads, with yellow-circled areas highlighting the regions that the heads focus on.
The feature prediction heads visualized in Figure 8 correspond to the indices 1, 5, 6, 17, and 18. These specific heads were selected to ensure a visually meaningful comparison, as some other heads produced fewer than six images that met the 0.75 threshold. The yellow highlighted areas indicate the regions that these heads attend to.
Figure 9 presents the same heads applied to a new batch of images, with a mask applied to remove areas that were not attended to.
Figure 9: Visualization of the same five feature prediction heads on a new batch of images, with non-attended regions masked out.
The model used for this visualization was pre-trained on the ImageNet-10 dataset with the following hyperparameters: batch size=256, epochs=1000, and learning rate=0.0003. All other settings were kept as defaults from our repository, where the pre-trained model files and complete experimental processes for all other experiments in this study are also available.
Appendix C Ablation Studies on Backbone Feature Extraction
This section provides additional ablation studies focused on evaluating the backbone feature extraction capabilities of our proposed framework. These experiments assess the contribution of various framework components by examining their impact on the backbone’s ability to extract meaningful features. Specifically, we analyze the performance of the backbone under different configurations by removing or modifying key components such as the instance projector, feature predictor, and data augmentation strategies.
The experiments are conducted on the STL-10 and ImageNet-10 datasets with the same batch size of 128, as used in the main experiments. We report the results using metrics such as Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), and Accuracy (ACC). The backbone performance is compared across several configurations to highlight the individual and combined impact of the modules on feature extraction.
Table 8: Effect of cosine decay schedule (Sch.) and gradient clipping (Clip.) on feature extraction.
Dataset Scheduler & Backbone
Grad Clipping NMI ARI ACC
STL-10 - $\mathbf{0.677\pm 0.020}$ $\mathbf{0.547\pm 0.031}$ $\mathbf{0.705\pm 0.039}$
$0.674\pm 0.015$ $0.545\pm 0.032$ $0.701\pm 0.034$
ImageNet-10 - $0.857\pm 0.021$ $0.772\pm 0.068$ $0.858\pm 0.071$
$\mathbf{0.886\pm 0.008}$ $\mathbf{0.844\pm 0.022}$ $\mathbf{0.918\pm 0.026}$
Table 9: Effect of feature-level head on feature extraction.
Dataset Feature-Level Backbone
NMI ARI ACC
STL-10 - $0.659\pm 0.016$ $0.530\pm 0.028$ $0.692\pm 0.030$
$\mathbf{0.674\pm 0.015}$ $\mathbf{0.545\pm 0.032}$ $\mathbf{0.701\pm 0.034}$
ImageNet-10 - $0.870\pm 0.022$ $0.809\pm 0.064$ $0.886\pm 0.058$
$\mathbf{0.886\pm 0.008}$ $\mathbf{0.844\pm 0.022}$ $\mathbf{0.918\pm 0.026}$
Table 10: Effect of single-view vs. dual-view data augmentation on feature extraction.
Dataset Augmentation Backbone
x1 x2 NMI ARI ACC
STL-10 - $0.660\pm 0.011$ $0.520\pm 0.022$ $0.683\pm 0.022$
$\mathbf{0.671\pm 0.019}$ $\mathbf{0.545\pm 0.032}$ $\mathbf{0.701\pm 0.034}$
ImageNet-10 - $\mathbf{0.889\pm 0.010}$ $\mathbf{0.851\pm 0.024}$ $\mathbf{0.923\pm 0.018}$
$0.886\pm 0.008$ $0.844\pm 0.022$ $0.918\pm 0.026$